Question 41 of 55beginner🔧 ApplyShort Answer3 marks

A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine (i) P(not A), (ii) P(not B) and (iii) P(A or B)

(i)
Answer

P(not A) = 0.58

(ii)
Answer

P(not B) = 0.52

(iii)
Answer

P(A or B) = 0.74

Explanation

This question tests the understanding of basic probability formulas. For parts (i) and (ii), we use the complement rule P(not A) = 1 - P(A). For part (iii), we apply the addition theorem of probability for two events that are not mutually exclusive.